OpenStudy (anonymous):

find the f-1 the inverse for f(x)= (x-1)^3 Help

OpenStudy (lukecrayonz):

switch y and x around, solve.

OpenStudy (anonymous):

y=(x-1)^3 x=(y-1)^3 than what do i do after that

myininaya (myininaya):

Solve for y.

OpenStudy (anonymous):

details please

terenzreignz (terenzreignz):

Well, best get rid of that ^3. How do you go about that?

OpenStudy (anonymous):

square root???

terenzreignz (terenzreignz):

almost... but that's for getting rid of ^2 (squares) How do you get rid of the ^3 (cube) ? ;)

OpenStudy (anonymous):

i thought thats where u come in place and tell me

terenzreignz (terenzreignz):

the cube root?

OpenStudy (anonymous):

i dont know

terenzreignz (terenzreignz):

Get the cube root of both sides of \[\Large x = (y-1)^3\]

terenzreignz (terenzreignz):

?? \[\Large \sqrt[3]x = \sqrt[3]{(y-1)^3 }\]

OpenStudy (anonymous):

than it x=\[\sqrt[3]{y-1?}\]

terenzreignz (terenzreignz):

Not exactly, no.

OpenStudy (anonymous):

im confused

terenzreignz (terenzreignz):

\[\Large \sqrt[3]{p^3}=\color{red}?\]

OpenStudy (anonymous):

x

OpenStudy (anonymous):

lol i dont know

terenzreignz (terenzreignz):

pardon? lol

OpenStudy (anonymous):

its been almost 10yrs ago since i have takin algebra in high school so i need major help as u see

terenzreignz (terenzreignz):

Shame, that. Perhaps you should recall your laws of exponents?

OpenStudy (anonymous):

ayy ayy captain